In many countries such as the US, when a pharmaceutical company discovers a new drug, it can apply to the government for a patent on the new drug. The patent gives the company the exclusive right to sell the new drug for a long period of time, such as 20 years. In other words, the pharmaceutical company is a monopolist in the market for the new drug.

Suppose the market demand for the new drug is shown as below:

Price

Quantity demanded

$100

0

90

100

80

200

70

300

60

400

50

500

40

600

30

700

20

800

10

900

0

1,000

  1. Suppose the marginal cost of producing the drug is a constant $10 per unit. What quantity would the profit-maximizing pharmaceutical company choose? What price would it charge?                                                          (6 marks)

Answer:

  1. When the patent runs out, other pharmaceutical companies quickly enter and begin selling so-called generic drugs that are chemically identical to the former monopolist’s brand-name drug. Suppose consumers are convinced that the former monopolist’s brand-name drug and other companies’ drugs are identical and the drug market becomes perfectly competitive.
  1. In terms of the characteristics of market structure, name any two differences between perfect competition and monopoly. What is in common for the two types of firms in maximizing profit?                                                      (4 marks)

Answer:

  1. If the market demand for drugs and marginal cost of producing drugs remain unchanged, what will be the new price and quantity of drug after the market structure changed?                          In what sense does this change improve the well-being of the society?        (10 marks)                                 
Answer #1

a) The company has invented a drug and received the patent for that. It means it will have a monopoly in the market for the period of patent. A monopoly can maximize the profit at the point where its marginal cost is equal to the marginal revenue. Its marginal cost is given as $10.

At price of $90 demand will be 100 so the revenue will be $9000.
MR = (9000 - 0) / (100 - 0) = 90

At price of $80 demand will be 200 and so that revenue will be $16000
MR = (16000 - 9000) / (200 - 100) = 70

We can create a table here.

Price Demand Revenue MR Profit
100 0 0.00 0
90 100 9000.00 90 8000
80 200 16000.00 70 14000
70 300 21000.00 50 18000
60 400 24000.00 30 20000
50 500 25000.00 10 20000
40 600 24000.00 -10 18000
30 700 21000.00 -30 14000
20 800 16000.00 -50 8000
10 900 9000.00 -70 0
0 1000 0.00 -90 -10000

We can observe that marginal cost of $10 is equal to marginal revenue at the price of $50 and quantity of 500.
Profit = Revenue - Cost
(50 * 500) - (10 * 50) = 20000

b) The monopoly and the perfect competition are the types of market. Here are few characteristics of each type.

Monopoly -
There is a single firm in the market which control the market.
There are barriers to enter and exit the market.
The company is price maker.
There is an imperfect knowledge about the market condition.

Perfect Competition -
There are a large number of identical firms in the market and nobody can control the market.
There are no barriers to enter of exit the market.
The firms are price taker and single company can not influence the price.
There is a perfect knowledge about the market condition.

However, firms in both the market seeks profit maximization and it can be achieved at the point where marginal revenue is equal to the marginal cost.

c) The expiry of the patent abolishes the exclusive right of the patent holding firm and other companies can also sell the identical drug in the market. If there are a lot of firms which enter the market then the market type could change to the perfect competition. In the perfect competition the firm are price taker and the price is equal to marginal cost.
Price = Marginal Cost
This means the market price will be $10 and quantity at that price will be 900.

The perfect competition has lowered the price while increasing the available quantity. The overall scenario increases welfare in the society from earlier situation of higher price and lower production. The perfcte competiton tends to be beneficial to the society.